# TraCS: Trajectory Collection in Continuous Space under Local Differential Privacy

### v1

@@ 2       $\tau_{i}=(a_{i},b_{i}),\tau_{i+1}=(a_{i+1},b_{i+1})$;
       $\varphi^{\prime}\leftarrow\mathcal{M}_{\circ}(\varphi;\varepsilon_{d})$;
@@ 5       $\mathcal{T}^{\prime}\leftarrow\mathcal{T}^{\prime}\cup\tau_{i+1}^{\prime}$;
6
7return $\mathcal{T}^{\prime}$;
@@
3.2.3. Workflow of TraCS-D
Combining the direction perturbation and distance perturbation, TraCS-D perturbs each private location in the trajectory.
Algorithm 1 presents the workflow of TraCS-D. It takes the location space $\mathcal{S}$, the private trajectory $\mathcal{T}$, and the privacy parameter $\varepsilon$ as input. Since TraCS-D relies on a non-private reference location, we add a dummy location $\tau_{0}^{\prime}$ to $\mathcal{T}$, which can be the starting coordinate of $\mathcal{S}$ or a randomly drawn location. Then, TraCS-D iterates over each pair of consecutive locations $\tau_{i}$ and $\tau_{i+1}$. For each private location $\tau_{i+1}$, it perturbs the direction (red block), and normalized distance (blue block) with privacy parameters $\varepsilon_{d}$ and $\varepsilon-\varepsilon_{d}$, respectively. The perturbed distance is then mapped along the perturbed direction (line 9-10). Line 12 updates the reference location to the perturbed location $\tau_{i+1}^{\prime}$ and performs the next iteration. Finally, it outputs the perturbed trajectory $\mathcal{T}^{\prime}$.
Algorithm 1 uses each $\tau_{i}^{\prime}$ as the reference location for the next perturbation. In fact, the reference location can be any non-private location in $\mathcal{S}$. We choose different $\tau_{i}^{\prime}$ for the consideration of alleviating the impact of a specific reference location.
3.2.4. Analysis of TraCS-D
@@
(Sketch) The direction perturbation mechanism $\mathcal{M}_{\circ}$ satisfies $\varepsilon_{d}$-LDP by its definition. For the distance perturbation, the randomness comes entirely from mechanism $\mathcal{M}_{-}$, which satisfies $(\varepsilon-\varepsilon_{d})$-LDP. The post-processing property preserves the same privacy level after linearly mapping the perturbed distance. Then each location satisfies $\varepsilon$-LDP due to Composition Theorem 2.2. Hence, the entire trajectory perturbed by TraCS-D satisfies $n\varepsilon$-LDP. Appendix A.1 provides the detailed proof. ∎
Complexity. TraCS-D (Algorithm 1) has $\Theta(n)$ time complexity and $\Theta(n)$ space complexity, where $n$ is the length of the trajectory.  The time complexity is $\Theta(n)$ because each line of Algorithm 1 is $\Theta(1)$. The space complexity is $\Theta(n)$ because it needs to store the perturbed trajectory $\mathcal{T}^{\prime}$.
The $\Theta(n)$ time complexity is optimal for per-location perturbation mechanisms. Among existing LDP mechanisms for discrete location spaces, even if ignore the complexity of the score function and sampling in the Exponential mechanism (Nissim et al., 2012), NGram (Cunningham et al., 2021b) has $\Theta(tn)$ time complexity, where $t$ is the number of predefined “gram”. It also includes a large constant factor due to its search to satisfy reachability constraints. ATP (Zhang et al., 2023) is $\Theta(mn)$, where $m$ is the size of the location space, due to its necessary trajectory merging step. LDPTrace (Du et al., 2023) has $\Theta(mn)$ time complexity, where $m$ is the number of grids; it also has a large constant factor due to the model training procedure on a set of trajectories.
Note that the $\Theta(n)$ space complexity is for the convenience of presenting the algorithm. It can be reduced to $\Theta(1)$ by updating each perturbed location $\tau_{i+1}^{\prime}$ in place within $\mathcal{T}$. This results in a lightweight memory footprint, which is crucial for mobile devices and edge computing.
Theorem 3.5.
@@
(Sketch) The MSE can be calculated by the closed-form expression of $\mathcal{M}_{\circ}$ and $\mathcal{M}_{-}$, allowing us to prove the convergence rate. Appendix A.2 provides the detailed proof. ∎
This theorem indicates that both the error of the perturbed direction and distance are reduced exponentially with the privacy parameter $\varepsilon$.  Note that the size of the distance space also affects the variance. Denote the output of $\mathcal{M}_{-}$ as a random variable $Y$ and the private input as $x$. After the linear mapping, the perturbed distance is $Y\cdot|\mathcal{D}_{r}|$ and the private distance is $x\cdot|\mathcal{D}_{r}|$, where $|\mathcal{D}_{r}|$ is the size of the distance space. Thus, the MSE of the perturbed distance is $\text{MSE}[Y\cdot|\mathcal{D}_{r}|]=|\mathcal{D}_{r}|^{2}\cdot\text{MSE}[Y]$. Therefore, a smaller $|\mathcal{D}_{r}|$ leads to a smaller MSE of the perturbed distance.
3.3. TraCS-C
TraCS-C decomposes the rectangular location space $\mathcal{S}$ into two independent distance spaces $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$, as shown in Figure 1(b).  Specifically, we can fix $(a_{\text{sta}},b_{\text{sta}})$ as the reference location, with longitude $a_{\text{sta}}$ as the $a$-axis and latitude $b_{\text{sta}}$ as the $b$-axis. Then any location $\tau_{i}\in\mathcal{S}$ can be represented as a pair of Cartesian coordinates $(d_{a},d_{b})$, where $d_{a}\in\mathcal{D}_{a}$ and $d_{b}\in\mathcal{D}_{b}$ are the distances from $\tau_{i}$ to the $a$-axis and $b$-axis, respectively. This representation is unique and reversible: any pair of distances $(d_{a},d_{b})$ can be mapped back to a unique location $\tau_{i}\in\mathcal{S}$.  Thus, if $\tau_{i}$ is a private location, an LDP mechanism applied to $\tau_{i}$ is essentially applied to $d_{a}$ and $d_{b}$. By Composition Theorem 2.2, we can achieve LDP for $\tau_{j}$ by applying perturbation mechanisms for $d_{a}$ and $d_{b}$, respectively.
3.3.1. Coordinates Perturbation
@@
Following the same idea of distance perturbation in TraCS-D, we normalize the distance $d_{a}$ and $d_{b}$ to $[0,1)$ by dividing $|\mathcal{D}_{a}|$ and $|\mathcal{D}_{b}|$, respectively. Then the mechanism $\mathcal{M}_{-}$ in Definition 3.3 provides $\varepsilon$-LDP for the normalized distances, and the linear mappings back to $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$ preserve the privacy guarantee.
Input: Location space $\mathcal{S}$, private trajectory $\mathcal{T}=\{\tau_{1},\tau_{2},\dots,\tau_{n}\}$, privacy parameter $\varepsilon$
@@ 3       $\tau_{i}=(a_{i},b_{i})$;
       $(\overline{d}_{a},\overline{d}_{b})\leftarrow(\frac{a_{i}-a_{\text{sta}}}{a_{\text{end}}-a_{\text{sta}}},\frac{b_{i}-b_{\text{sta}}}{b_{\text{end}}-b_{\text{sta}}})$;
@@ 4       $\overline{d}^{\prime}_{a}\leftarrow\mathcal{M}_{-}(\overline{d}_{a};\varepsilon/2)$;
       $\overline{d}^{\prime}_{b}\leftarrow\mathcal{M}_{-}(\overline{d}_{b};\varepsilon/2)$;
@@
       $(d_{a}^{\prime},d_{b}^{\prime})\leftarrow(\overline{d}^{\prime}_{a}|\mathcal{D}_{a}|,\overline{d}^{\prime}_{b}|\mathcal{D}_{b}|)$;
[…]
5       $\tau^{\prime}_{i}=(a_{\text{sta}},b_{\text{sta}})+(d_{a}^{\prime},d_{b}^{\prime})$;
@@ 6       $\mathcal{T}^{\prime}\leftarrow\mathcal{T}^{\prime}\cup\{\tau^{\prime}_{i}\}$;
7
8return $\mathcal{T}^{\prime}$;
@@
Algorithm 2 shows the workflow of TraCS-C. It takes the location space $\mathcal{S}$, the private trajectory $\mathcal{T}$, and the privacy parameter $\varepsilon$ as input, and outputs the perturbed trajectory $\mathcal{T}^{\prime}$. For each location $\tau_{i}=(a_{i},b_{i})$ in $\mathcal{T}$, Line 4 computes its coordinates $(d_{a},d_{b})$ and normalizes them to $(\overline{d}_{a},\overline{d}_{b})\in[0,1)\times[0,1)$. Lines 5-6 perturb the normalized coordinates $\overline{d}_{a}$ and $\overline{d}_{b}$ using $\mathcal{M}_{-}$ with privacy parameter $\varepsilon/2$. Then Line 7 maps the perturbed normalized coordinates back to $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$, and Line 8 converts them to the longitude-latitude representation $\tau^{\prime}_{i}$.
3.3.2. Analysis of TraCS-C
Theorem 3.6.
@@ 3.3.2. Analysis of TraCS-C
(Sketch) Algorithm 2 uses mechanism $\mathcal{M}_{-}$ twice, each with a privacy parameter $\varepsilon/2$. By Composition Theorem 2.2, their composition satisfies $\varepsilon$-LDP. The subsequent linear mappings to $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$ are post-processing steps that preserve the same privacy. Hence, each perturbed location satisfies $\varepsilon$-LDP, and the whole trajectory satisfies $n\varepsilon$-LDP. Appendix A.3 provides the details. ∎
Complexity. TraCS-C (Algorithm 2) has $\Theta(n)$ time complexity and $\Theta(n)$ space complexity, where $n$ is the length of the trajectory.  The time complexity is $\Theta(n)$ because each line of Algorithm 2 is $\Theta(1)$. The space complexity is $\Theta(n)$ because it needs to store the perturbed trajectory $\mathcal{T}^{\prime}$. Note that the $\Theta(n)$ space complexity is for the convenience of presenting the algorithm. It can also be reduced to $\Theta(1)$ by updating each perturbed location $\tau_{i}^{\prime}$ in place within $\mathcal{T}$.
3.4. Rounding to Discrete Space
@@
If the location space is a set of location points, we can round the perturbed location to the nearest location point. Formally, assume $\mathcal{P}=\{p_{1},\ldots,p_{m}\}\subseteq\mathcal{S}$ is a set of location points, and $\tau_{i}^{\prime}$ is the perturbed location of $\tau_{i}$ in TraCS. Then, the discretized TraCS outputs $\operatorname*{arg\,min}_{p_{j}\in\mathcal{P}}||\tau_{i}^{\prime}-p_{j}||_{2}$. The time complexity of rounding to the nearest location point is $\mathcal{O}(m)$ in the worst case. In practice, this time complexity can be significantly reduced by using heuristic or approximation algorithms.
3.5. Discussion and Extensions
This subsection discusses the comparison between TraCS-D and TraCS-C, and their extensions to other types of location spaces.
3.5.1. Privacy Comparison Between TraCS-D and TraCS-C
Although TraCS-D and TraCS-C provide the same level of privacy guarantee, they have different privacy interpretations.
In TraCS-D, the LDP guarantee is provided for the direction information and the subsequent distance information. This means that when the usage of TraCS-D and the perturbed trajectory $\mathcal{T}^{\prime}$ is publicly known, any observer can hardly infer the private direction $\varphi:\tau^{\prime}_{i}\to\tau_{i+1}$ and the private distance $r(\varphi):|\tau^{\prime}_{i}\to\tau_{i+1}|$ from the known $\tau^{\prime}_{i}$ and $\tau^{\prime}_{i+1}$.  Although the distance space $\mathcal{D}_{r(\varphi)}$ relies on a specific direction, it does not leak the direction information, as the direction space $\mathcal{D}_{\varphi}=[0,2\pi)$ is independent of any location and TraCS-D uses the perturbed direction, which preserves LDP.
In TraCS-C, the LDP guarantee is provided for the two independent distance spaces $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$. This means that when the usage of TraCS-C and the perturbed trajectory $\mathcal{T}^{\prime}$ is publicly known, any observer can hardly infer the private distances $d_{a}$ and $d_{b}$.
3.5.2. Impact of the Size of $\mathcal{S}$

### v2

@@ 2  $\tau_{i}=(a_{i},b_{i}),\tau_{i+1}=(a_{i+1},b_{i+1})$;
3  $\varphi\leftarrow\text{atan2}(b_{i+1}-b_{i},a_{i+1}-a_{i})$;
@@
4  $\varphi^{\prime}\leftarrow\mathcal{M}_{\circ}(\varphi;\varepsilon_{d})$;
@@ 7  $\mathcal{T}^{\prime}\leftarrow\mathcal{T}^{\prime}\cup\tau_{i+1}^{\prime}$;
8
9return $\mathcal{T}^{\prime}$;
@@
3.2.4 Analysis of TraCS-D
@@
(Sketch) The direction perturbation mechanism $\mathcal{M}_{\circ}$ satisfies $\varepsilon_{d}$-LDP by its definition. For the distance perturbation, the randomness comes entirely from mechanism $\mathcal{M}_{-}$, which satisfies $(\varepsilon-\varepsilon_{d})$-LDP. The post-processing property preserves the same privacy level after linearly mapping the perturbed distance. Then each location satisfies $\varepsilon$-LDP by Composition Theorem 1 (or by computing the 2D pdfpdf ratio in the LDP definition). Hence, the entire perturbed trajectory satisfies $n\varepsilon$-LDP. Appendix B.2 provides details. ∎
Complexity. TraCS-D (Algorithm 1) has $\Theta(1)$ time complexity for each location, as each line is $\Theta(1)$. Thus, the entire algorithm has $\Theta(n)$ time complexity, where $n$ is the length of the trajectory. The space complexity is also $\Theta(n)$ because it needs to store the perturbed trajectory $\mathcal{T}^{\prime}$.
The $\Theta(n)$ time complexity is optimal for per-location perturbation mechanisms. Among existing LDP mechanisms for discrete location spaces, even if we ignore the complexity of the score function and sampling in the Exponential mechanism [27], NGram [9] has $\Theta(tn)$ time complexity, where $t$ is the number of predefined “gram”. It also includes a large constant factor due to its search to satisfy reachability constraints. L-SRR [29] has $\Theta(mn)$ time complexity, where $m$ is the number of locations, due to the grouping of locations. ATP [37] also has $\Theta(mn)$ time complexity, because of its trajectory merging step.
Note that the $\Theta(n)$ space complexity is for the convenience of presenting the algorithm. It can be reduced to $\Theta(1)$ by only updating each perturbed location $\tau_{i+1}^{\prime}$ in place within $\mathcal{T}$. This results in a lightweight memory footprint, which is crucial for edge computing.
Theorem 3.
@@
(Sketch) The MSE can be calculated by the closed-form expression of $\mathcal{M}_{\circ}$ and $\mathcal{M}_{-}$, allowing us to prove the convergence rate. Appendix B.3 provides the details. ∎
This theorem indicates that both the error of the perturbed direction and distance are reduced exponentially with the privacy parameter $\varepsilon$. Note that the size of the distance space also affects the variance. Specifically, denote the output of $\mathcal{M}_{-}$ as a random variable $Y$ and the sensitive input as $x$. After the linear mapping, the perturbed distance is $Y\cdot|\mathcal{D}_{r}|$ and the sensitive distance is $x\cdot|\mathcal{D}_{r}|$, where $|\mathcal{D}_{r}|$ is the size of the distance space. Thus, the MSE of the perturbed distance is $\text{MSE}[Y\cdot|\mathcal{D}_{r}|]=|\mathcal{D}_{r}|^{2}\cdot\text{MSE}[Y]$. Therefore, a smaller $|\mathcal{D}_{r}|$ leads to a smaller MSE of the perturbed distance.
3.3 TraCS-C
TraCS-C decomposes the rectangular location space $\mathcal{S}$ into two independent distance spaces $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$, as shown in Figure 1(b). Specifically, we can fix $(a_{\text{sta}},b_{\text{sta}})$ as the reference location, with longitude $a_{\text{sta}}$ as the $a$-axis and latitude $b_{\text{sta}}$ as the $b$-axis. Then any location $\tau_{i}\in\mathcal{S}$ can be represented as a pair of Cartesian coordinates $(d_{a},d_{b})$, where $d_{a}\in\mathcal{D}_{a}$ and $d_{b}\in\mathcal{D}_{b}$ are the distances from $\tau_{i}$ to the $a$-axis and $b$-axis, respectively. This representation is unique and reversible: any pair of distances $(d_{a},d_{b})$ can be mapped back to a unique location $\tau_{i}\in\mathcal{S}$. Thus, if $\tau_{i}$ is a sensitive location, an LDP mechanism applied to $\tau_{i}$ is essentially applied to $d_{a}$ and $d_{b}$. By Composition Theorem 1, we can achieve LDP for $\tau_{j}$ by applying perturbation mechanisms for $d_{a}$ and $d_{b}$, respectively.
3.3.1 Coordinates Perturbation
@@
Following the same idea of distance perturbation in TraCS-D, we normalize the distance $d_{a}$ and $d_{b}$ to $[0,1)$ by dividing $|\mathcal{D}_{a}|$ and $|\mathcal{D}_{b}|$, respectively. Then the mechanism $\mathcal{M}_{-}$ in Definition 6 provides $\varepsilon$-LDP for the normalized distances, and the linear mappings back to $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$ preserve the privacy guarantee.
Input: Rectangular location space $\mathcal{S}$, sensitive trajectory $\mathcal{T}=\{\tau_{1},\tau_{2},\dots,\tau_{n}\}$, privacy parameter $\varepsilon$
@@ 3  $\tau_{i}=(a_{i},b_{i})$;
[…]
4  $(\overline{d}_{a},\overline{d}_{b})\leftarrow(\frac{a_{i}-a_{\text{sta}}}{a_{\text{end}}-a_{\text{sta}}},\frac{b_{i}-b_{\text{sta}}}{b_{\text{end}}-b_{\text{sta}}})$;
@@ 5  $\overline{d}^{\prime}_{a}\leftarrow\mathcal{M}_{-}(\overline{d}_{a};\varepsilon/2)$;
6  $\overline{d}^{\prime}_{b}\leftarrow\mathcal{M}_{-}(\overline{d}_{b};\varepsilon/2)$;
@@
7  $(d_{a}^{\prime},d_{b}^{\prime})\leftarrow(\overline{d}^{\prime}_{a}|\mathcal{D}_{a}|,\overline{d}^{\prime}_{b}|\mathcal{D}_{b}|)$, $\tau^{\prime}_{i}=(a_{\text{sta}},b_{\text{sta}})+(d_{a}^{\prime},d_{b}^{\prime})$;
@@ 8  $\mathcal{T}^{\prime}\leftarrow\mathcal{T}^{\prime}\cup\{\tau^{\prime}_{i}\}$;
9
10return $\mathcal{T}^{\prime}$;
@@
Algorithm 2 shows the workflow of TraCS-C. It takes the rectangular location space $\mathcal{S}$, the sensitive trajectory $\mathcal{T}$, and the privacy parameter $\varepsilon$ as input, and outputs the perturbed trajectory $\mathcal{T}^{\prime}$. For each location $\tau_{i}=(a_{i},b_{i})$ in $\mathcal{T}$, Line 4 computes its coordinates $(d_{a},d_{b})$ and normalizes them to $(\overline{d}_{a},\overline{d}_{b})\in[0,1)\times[0,1)$. Lines 5-6 perturb the normalized coordinates $\overline{d}_{a}$ and $\overline{d}_{b}$ using $\mathcal{M}_{-}$ with privacy parameter $\varepsilon/2$. Then Line 8 maps the perturbed normalized coordinates back to $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$, and converts them to the longitude-latitude representation $\tau^{\prime}_{i}$.
3.3.2 Analysis of TraCS-C
Privacy and complexity of TraCS-C can be analyzed similarly to TraCS-D.
Theorem 4.
@@ 3.3.2 Analysis of TraCS-C
(Sketch) Algorithm 2 uses mechanism $\mathcal{M}_{-}$ twice, each with a privacy parameter $\varepsilon/2$. By Composition Theorem 1 (or by computing the 2D pdfpdf ratio in the LDP definition), their composition satisfies $\varepsilon$-LDP. The subsequent linear mappings to $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$ are post-processing steps that preserve the same privacy. Hence, each perturbed location satisfies $\varepsilon$-LDP, and the whole trajectory satisfies $n\varepsilon$-LDP. Appendix B.4 provides details. ∎
Complexity. TraCS-C (Algorithm 2) has $\Theta(n)$ time complexity and $\Theta(n)$ space complexity, where $n$ is the length of the trajectory. The time complexity is $\Theta(n)$ because each line of Algorithm 2 is $\Theta(1)$. The space complexity is $\Theta(n)$ because it needs to store the perturbed trajectory $\mathcal{T}^{\prime}$. Note that the $\Theta(n)$ space complexity is for the convenience of presenting the algorithm. It can also be reduced to $\Theta(1)$ by updating each perturbed location $\tau_{i}^{\prime}$ in place within $\mathcal{T}$, which benefits edge devices with limited computation resources.
3.4 Rounding to Discrete Space
@@
[…]
Heuristic reference location. The possible locations in a discrete space are generally public knowledge, which aids in heuristically selecting a better reference location for TraCS-D. The key observation is that a smaller distance space $\mathcal{D}_{r(\varphi)}$ leads to a smaller distance perturbation error. Therefore, it is beneficial to select a reference location that minimizes the size of $\mathcal{D}_{r(\varphi)}$ to other locations.
Specifically, the direction perturbation from the reference location is likely to be within the dominant sector shown in Figure 2. If a dominant sector contains most of the locations (resulting in smaller rounding error) and is near the boundary of the location space (resulting in smaller distance space), then the corresponding reference location is a good choice.
3.5 Discussions and Extensions
This subsection discusses the comparison between TraCS-D and TraCS-C, including their privacy, sensitivity to the location space, and extensions to other types of location spaces.
3.5.1 Comparison of Privacy Between TraCS-D and TraCS-C
Although TraCS-D and TraCS-C provide the same level of privacy quantified by $\varepsilon$-LDP—more specifically, location-level (or event-level) LDP as described in surveys [26, 3]—they have different privacy interpretations.
In TraCS-D, the LDP guarantee is provided for the direction information and the subsequent distance information. This means that when the usage of TraCS-D and the perturbed trajectory $\mathcal{T}^{\prime}$ is publicly known, any observer can hardly infer the sensitive direction $\varphi:\tau^{\prime}_{i}\to\tau_{i+1}$ and the sensitive distance $r(\varphi):|\tau^{\prime}_{i}\to\tau_{i+1}|$ from the known $\tau^{\prime}_{i}$ and $\tau^{\prime}_{i+1}$. Although the distance space $\mathcal{D}_{r(\varphi)}$ relies on a specific direction, it does not leak the direction information, as the direction space $\mathcal{D}_{\varphi}=[0,2\pi)$ is independent of any location and TraCS-D uses the perturbed direction.
In TraCS-C, the LDP guarantee is provided for the two independent distance spaces $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$. This means that when the usage of TraCS-C and the perturbed trajectory $\mathcal{T}^{\prime}$ is publicly known, any observer can hardly infer the sensitive distances $d_{a}$ and $d_{b}$.
3.5.2 Impact of the Size of $\mathcal{S}$